p= (z+qy)2 how to solve it by Jacobi 's method



1
Expert's answer
2021-05-04T13:07:55-0400

:Here the given partial differential equation is

"f(x, y, z, p, q) = p \u2212 (qy + z)^2 = 0, ~~~~~~~-(1)"


Now, the auxiliary equations are

"\dfrac{dp}{\frac{∂f}{∂x} +p \frac{∂f}{∂z}}=\dfrac{dp}{\frac{∂f}{∂y} + q\frac{∂f}{∂z}}=\dfrac{dz}{−p\frac{∂f}{∂p} − q\frac{∂f}{∂q}}=\dfrac{dx}{−\frac{∂f}{∂p}}=\dfrac{dy}{−\frac{∂f}{∂q}}"


or,

"\dfrac{dp}{2p(qy + z)}=\dfrac{dq}{2q(qy + z)}=\dfrac{dz}{(−p)(−1) − q.2(qy + z)y}=\dfrac{dx}{−(−1)} =\dfrac{dy}{−2y(qy + z)}"


Taking the 1st and 5th ratios, we get

"\dfrac{dp}{p}+\dfrac{dy}{y}= 0,"


Integrating, "log p + log y = log a" , which gives, "p =\dfrac{a}{y}."


Substitute in (1), we have,"\dfrac{a}{y} = (qy + z)^2" , which gives, "q =\sqrt{\dfrac{a}{y}}-\dfrac{z}{y}"


Substituting these value of p and q in "dz = p dx + q dy" , we get,


"dz =\dfrac{a}{y}dx + (\sqrt{\dfrac{a}{y}}−\dfrac{z}{y}) dy"


or

"y dz + z dy = a dx +\sqrt{\dfrac{a}{y}}dy"


After integration, we get, "yz = ax +2\sqrt{(ay)}+ b" , which is a complete integral

with a and b are arbitrary constants.